Fermi National Laboratory


Accelerator Update for January 9, 2006 - January 11, 2006

Monday January 9

The midnight shift began with Operations monitoring store #4586, the Recycler Ring (RR) stash at 114.8 E10, Pbar stack at 9.6 E10. MiniBooNE is off for mode change, NuMI and MCenter are off because of shots to the Recycler.

Pbar returns to stacking at 8:22 AM, beam also returns to NuMI and MCenter.

MiniBooNE came back on at 10:48 AM in the Horn Off running mode.

Operations put the Recycler into shot setup at 3:44 PM to stash antiprotons.

Pbar returns to stacking at 4:49 PM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 10:25 PM to stash antiprotons.

Pbar returns to stacking at 11:13 PM, beam also returns to MCenter.

Tuesday January 10

The midnight shift began with Operations monitoring store #4586, the Recycler Ring (RR) stash at 245.1 E10, Pbar stack at 27.4 E10, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the Recycler into shot setup at 3:37 AM to stash antiprotons.

Pbar returns to stacking at 4:28 AM, beam also returns to MCenter.

NuMI was turned off at 6:31 AM for a planned access.

Store #4586 was terminated at 8:22 AM.

Shot setup begins at 8:37AM.

Operations established store 4587 at 11:08 AM with an initial luminosity of 149.64 E30.

Operations put the Recycler into shot setup at 11:15 AM to stash antiprotons.

Pbar returns to stacking at 11:50 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 2:23 PM to stash antiprotons.

An internal coil water leak was found on the MCenter Analyzing Magnet MC7AN1 (Jolly Green Giant) at 2:25 PM. Experts are discussing repair options.

Pbar returns to stacking at 3:20 PM, MCenter off due to water leak on analyzing magnet.

NuMI access is complete, beam returns at 5:34 PM.

MCenter turns back on at 5:51 PM. They are currently running without their analyzing magnet MC7AN1.

Operations put the Recycler into shot setup at 8:15 PM to stash antiprotons.

Pbar returns to stacking at 11:18 PM, beam also returns to MCenter.

Wednesday January 11

The midnight shift began with Operations monitoring store #4587, the Recycler Ring (RR) stash at 132.5 E10, Pbar stack at 24.1 E10, and with MiniBooNE, NuMI, and MCenter taking beam.

An MCenter quad (MC6Q5) tripped off on high temperature at 2:02 AM. The magnet was reset and turned back on. This magnet tripped off two more times during the shift, the last being at 5:23 AM. The magnet was left off after this last trip.

Operations put the Recycler into shot setup at 3:54 AM to stash antiprotons.

Pbar returns to stacking at 4:39 AM.

MiniBooNE turned off for a scheduled access at 7:03 AM.

The Plan for Wednesday

Keep the store, stack, stash, run beam to NuMI. MiniBooNE is off for access, MCenter is off for a water leak on their analyzing magnet and for temperature problems on a quad magnet. End of store studies this afternoon.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/11/2006   email Fermilab